# Sampling with time-changed Markov processes

### v1

@@ Abstract.
We study time-changed Markov processes to speed up the convergence of Markov chain Monte Carlo (MCMC) algorithms in the context of multimodal distributions and rare event simulation. The time-changed process is defined by adjusting the speed of time of a base process via a user-chosen, state-dependent function. We apply this framework to several Markov processes from the MCMC literature, such as Langevin diffusions and piecewise deterministic Markov processes, obtaining novel modifications of classical algorithms and also re-discovering known MCMC algorithms. We prove theoretical properties of the time-changed process under suitable conditions on the base process, focusing on connecting the stationary distributions and qualitative convergence properties such as geometric and uniform ergodicity, as well as a functional central limit theorem. A comparison with the framework of space transformations is provided, clarifying the similarities between the approaches. Throughout the paper we give various visualisations and numerical simulations on simple tasks to gain intuition on the method and its performance.
@@ 1. Introduction
Markov chain Monte Carlo (MCMC) algorithms are the standard approach to obtain samples from a target distribution $\mu$ on a state space $\mathrm{E}\subset\mathbb{R}^{d}$. These methods are based on the simulation of a Markovian process with suitable properties ensuring that its law converges to $\mu$ as the physical time of the process grows. In this work we shall consider algorithms that rely on a continuous-time Markov processes $X$ that satisfies a law of large numbers (LLN), i.e.
$\frac{1}{T}\int_{0}^{T}f(X_{t})\textnormal{d}t\,\,\xrightarrow[a.s.]{T\rightarrow\infty}\int_{\mathrm{E}}f(x)\mu(\textnormal{d}x)$ (1)
for any observable $f$ in a suitable class (e.g. $\mu$-integrable functions). The focus of our paper shall be on improving the performance of MCMC algorithms in challenging settings that arise in practice, such as when $\mu$ is multimodal or when $\mu$ is heavy-tailed and we are interested in rare event estimation. In particular, multimodal distributions pose a serious challenge for MCMC algorithms because the process is attracted by the mode it finds itself in, to the extent that it struggles to hop to another mode. This phenomenon can be understood applying (1) to the observable $f(x)=\mathbbm{1}_{A}(x)$, where $A$ is a measurable set that the process must traverse in order to reach a different mode. The LLN gives that the fraction of time $X$ spends in $A$ converges to $\mu(A):=\int_{A}\mathbbm{1}_{A}(x)\mu(\textnormal{d}x)$, and hence it is clear that $X$ can only visit $A$ rarely when $\mu(A)$ is small. As a consequence, a large computational effort is required before the process leaves the neighbourhood of the mode it is exploring, since this means going through (and thus spending time) in a region of low density.
@@
In this paper, we introduce and discuss the framework of sampling with time-changed Markov processes to obtain more efficient MCMC algorithms in difficult tasks as the ones described above. The fundamental idea is to apply a random time-change to a base Markov process in such a way that time accelerates in regions that we wish the process visits often, but where, with large probability, it cannot spend too much time without breaking the law of large numbers. The interpretation of these regions depends on the task. For instance, these are the low-density areas that separate modes in the multimodal setting, or the tails of the target distribution in the case of rare events estimation. In practice, the time-change is regulated by a speed function, $s$, which maps the state space to strictly positive real numbers and accelerates or decelerates time based on the current state of the process. More formally, we introduce a Markov process $X$ by time-changing a base process, $Y$, as follows:
$X_{t}=Y_{r(t)},\quad\text{where }\,r(t)=\int_{0}^{t}s(X_{u})\,\textnormal{d}u.$ (2)
Importantly, we establish that $X$ has $\mu$ as its stationary distribution when $Y$ is stationary with respect to the probability distribution $\tilde{\mu}$, which is proportional to $s\mu$. This property reveals a connection between time-changed processes and importance sampling. In importance sampling, samples from $\tilde{\mu}$ are re-weighted by $\frac{\mu}{\tilde{\mu}}$ to approximate expectations with respect to $\mu$. Similarly, the paths of the time-changed process $X$ are obtained by re-weighting paths from the $\tilde{\mu}$-stationary base process. Here, the time change assigns importance weights, where large weights are assigned to states where the speed function is small, that is where the time-changed process moves slowly. The speed function, defined as $\frac{\tilde{\mu}}{\mu}$ up to a constant factor, can then be interpreted as the reciprocal of the importance weight. Intuitively, a well-chosen speed function defines a distribution $\tilde{\mu}$ that makes the task of interest easier to solve, for example connecting the modes of $\mu$ when it is multimodal. As in importance sampling, it is essential to choose speed functions bounded below by a strictly positive constant, ensuring that $\tilde{\mu}$ has heavier tails than $\mu$.
@@
In this work we assume the base process, $Y$, is a well studied Markov process and give conditions on $Y$ and $s$ that ensure desirable properties of the time-changed process $X$, such as the LLN, qualitative convergence rates to the stationary distribution, and functional central limit theorems. In particular, suitable choices of $s$ ensure that the convergence of $X$ is exponential in time even if $Y$ converges only polynomially or less. Moreover, we provide conditions for uniform ergodicity of $X$, that is exponential convergence of the law to the target independently of the initial condition. Throughout the paper we highlight the numerous connections to the existing literature and apply our framework to several Markov processes that are widely used in the context of MCMC algorithms.
@@
  • Section 5 discusses the relation between time changes and space transformations, illustrating the similarities and differences of the two methods. In particular, we shall observe how the speed function in our framework plays a similar role as the Jacobian determinant of the diffeomorphism that defines the space transformations. We analyse this connection considering a novel diffeomorphism, which maps $\mathbb{R}^{d}$ to points inside the $d$-dimensional unit sphere.
@@
Related work
To the best of our knowledge, the idea of sampling with a time-change of a Markov process was first discussed by Roberts and Stramer (2002a). The authors consider a time-changed overdamped Langevin diffusion corresponding to speed functions of the particular form $s(x)=\mu^{-\alpha}(x)$ for $\alpha\in(0,1)$. The focus of Roberts and Stramer (2002a) is then limited to a specific choice of process and speed function, while in this paper we define and study the framework in generality. As mentioned above, the time-change of the ZZP was already considered by Vasdekis and Roberts (2023), but the process was introduced with ad-hoc arguments and without indicating this connection. The approach of time-change is also deeply connected to importance sampling (Robert and Casella, 2004; Owen, 2013). Indeed, computing ergodic averages with respect to a time-changed process is equivalent to obtaining re-weighted ergodic averages of the base process for a random time horizon (see Equation (10)). In this direction, Chak et al. (2023) studies the optimal choice of “proposal distribution” $\tilde{\mu}$ to use as target for the overdamped Langevin diffusion, then adjusting the obtained sample by importance weights. We shall revisit this connection in Example 3.8. It is worth mentioning that the speed function $s(x)=\mu^{-\alpha}(x)$ with $\alpha=1$ was shown to be optimal across all diffusion coefficients for the overdamped Langevin diffusion in a specific (multimodal) setting when the state space is included in a compact set (Lelièvre et al., 2024). In the discrete-time world, Andral et al. (2024) use as base process a Markov chain with stationary distribution $\tilde{\mu}$ and adjust it by letting it spend a random number of iterations at each state. This is closely connected to some of the ideas we discuss in Section 2.2. Similar ideas were introduced by Malefaki and Iliopoulos (2008) in a semi-Markov implementation. The frameworks of time changes and space transformations are similar when the main goal is to increase the frequency of visits to the tails, or equivalently to obtain a uniformly ergodic process. This aspect is discussed in Section 5, where we discuss particularly the connection to Johnson and Geyer (2012) and Yang et al. (2024). Finally, we mention that a first version of this work appeared in the PhD thesis of the first author (Bertazzi, 2023).
@@ 2.2. Estimating expectations with time-changed Markov processes
    In cases where simulating the process $X$ is challenging, we can estimate $\mu(f)$ by re-weighting paths of the base process $Y$. Indeed, the change of variables $\tilde{t}=r(t)$, with Jacobian $s(Y_{\tilde{t}})$, gives
    $\frac{1}{T}\int_{0}^{T}f(X_{t})\textnormal{d}t\,=\,\frac{1}{T}\int_{0}^{\tilde{T}}\frac{f(Y_{t})}{s(Y_{t})}\textnormal{d}t,$ (10)
    where $\tilde{T}$ is such that $r^{-1}(\tilde{T})=T$. Theorem 2.3 guarantees that the right hand side of (10) is a consistent estimator of $\mu(f)$. The estimator (10) requires the simulation of $Y$ for a random time $\tilde{T}$, but we can alternatively fix a deterministic time horizon $T$ and use the estimator
@@
    which is equivalent to simulating $X$ until time $r^{-1}(T)$. The estimator (11) converges to $\mu(f)$ since $T^{-1}r^{-1}(T)\to(\mu(s))^{-1}.$ We remark that for both (10) and (11) it is necessary to compute the inverse time change, $r^{-1}(T)$. The estimator (11) essentially coincides with doing self-normalised importance sampling for the base process $Y$, where the unnormalised weights are $(s(Y_{t}))^{-1}$. The estimator (11) is considered in Chak et al. (2023) for a particular choice of the base process, $Y$. Similarly to importance sampling, this approach allows the estimation of the normalising constant of $\mu$. Indeed, let $\mu(x)=Z^{-1}\exp(-\mathrm{U}(x))$, where $Z$ is unknown. We can estimate $Z$ with
    $\frac{r^{-1}(T)}{T\int_{\mathrm{E}}s(x)\exp(-\mathrm{U}(x))\textnormal{d}x}\,\,\xrightarrow[a.s.]{T\rightarrow\infty}\,\,Z,$ (12)
    which can be used when one has access to the quantity $\int_{\mathrm{E}}s(x)\exp(-\mathrm{U}(x))\textnormal{d}x$.
@@
    We now provide a third option in the case where directly simulating $X$ or the random time-change $r^{-1}(T)$ is not feasible. An alternative strategy is to use as base process a jump process with jump rate $\tilde{\lambda}(y)=1$ for all $y\in\mathrm{E}$, and jump kernel $\widetilde{Q}$, that has $\tilde{\mu}$ as stationary distribution. The kernel $\widetilde{Q}$ can be any kernel from the literature on discrete time MCMC algorithms, or also the transition kernel of a continuous time Markov process after some fixed time $\bar{t}$. In either case, the time changed process $X_{t}=Y_{r(t)}$ is a jump process with rate $\lambda(x)=s(x)$ and kernel $Q=\widetilde{Q}$ (see Theorem 3.1 below), with stationary distribution $\mu$. Fixing a time horizon $T$, we can directly simulate $X$ and obtain a skeleton chain $(X_{T_{n}},T_{n})_{n=0}^{N_{T}+1}$, where $T_{0}=0$ and $X_{0}=x$, $T_{n}$ is the $n$-th jumping event time, $N_{T}$ is the time of the last event before time $T$, while $T_{N_{T}+1}=T$ and $X_{t}=X_{N_{T}}$ for $t\in[T_{N_{T}},T_{N_{T}+1}]$. The construction of $X$ between any two jump times, say $T_{n}$ and $T_{n+1}$, is as follows. Conditionally on $(X_{T_{n}},T_{n})$, we first find the next jump time using that $T_{n+1}\sim T_{n}+\mathrm{Exp}(s(X_{T_{n}}))$. Then, we set $X_{t}=X_{T_{n}}$ for $t\in(T_{n},T_{n+1})$, and draw $X_{T_{n+1}}\sim\widetilde{Q}(X_{T_{n}},\cdot)$. This process is straightforward to simulate and has stationary distribution $\mu$ under the conditions of Theorem 2.3. This procedure suggests the following estimator
    $\frac{1}{T}\int_{0}^{T}f(X_{t})\textnormal{d}t\,=\,\frac{1}{T}\sum_{n=0}^{N_{T}}f(X_{T_{n}})(T_{n+1}-T_{n})\,.$ (13)
    One can “regularise” the estimator (13) by instead discretising the path $(X_{t})_{t\in[0,T]}$, that is using
@@ 3.1. Time-changes of piecewise deterministic Markov processes (PDMPs)
3.2. Running example: the Zig-Zag process
@@
3.2.1. One-dimensional bimodal target
@@ 3.2.2. Tail exploration
3.3. Other piecewise deterministic Markov processes
Taking advantage of Theorem 3.1, we now obtain several novel PDMPs with stationary distribution $\mu$. Using the same ideas as in Section 3.2, it is possible to define time-changes of PDMPs such as those designed for subsampling (see e.g. Bierkens et al. (2019a)), or of other PDMPs such as Wu and Robert (2020); Bierkens et al. (2020).
@@ 3.4. Diffusion processes
Assuming this SDE is well posed, we find that the solution has the measure with density $\mu(x,v)\propto\exp(-\mathrm{U}(x)-(1/2)\lvert v\rvert^{2})$ as stationary distribution.

### v2

@@ Abstract.
We study time-changed Markov processes to speed up the convergence of Markov chain Monte Carlo (MCMC) algorithms. The time-changed process is defined by adjusting the speed of time of a base process via a user-chosen, state-dependent function. We explore the properties of such transformations and apply this idea to several Markov processes from the MCMC literature, such as Langevin diffusions and piecewise deterministic Markov processes, obtaining novel modifications of classical algorithms and also re-discovering known MCMC algorithms. We prove theoretical properties of the time-changed process under suitable conditions on the base process, focusing on connecting the stationary distributions and qualitative convergence properties such as geometric and uniform ergodicity, as well as a functional central limit theorem. We also provide a comparison with the framework of space transformations, clarifying the similarities between the approaches. Throughout the paper we give various visualisations and numerical simulations on simple tasks to gain intuition on the method and its performance. Finally, we provide numerical simulations to gain intuition on the method and its performance on benchmark problems. Our results indicate a performance improvement in the context of multimodal distributions and rare event simulation.
@@ 1. Introduction
The need to generate samples from complex probability distributions is ubiquitous in Bayesian statistics and beyond (e.g. Brooks et al. (2011), Stuart (2010)). The standard technique for achieving this is Markov chain Monte Carlo (MCMC), in which a Markovian process is constructed such that its distribution converges to the target distribution as the process evolves. The outcomes of the process are then used as samples from the distribution of interest. While the original MCMC algorithms used discrete-time processes, continuous-time processes have since gained popularity. These include diffusion processes (Roberts and Tweedie, 1996), Markov jump processes (Green, 1995; Power and Goldman, 2019), and piecewise deterministic Markov processes (PDMPs) (Fearnhead et al., 2018). However, many of these processes struggle with targets commonly encountered in practical settings, such as multi-modal or heavy tailed distributions. For example, in the case of multi-modal targets, the processes tend to be attracted to the mode they find themselves in, to the extent that they struggle to hop to another mode. In the case of heavy-tailed distributions, they fail to explore the tails of the distribution in a systematic and consistent way.
@@
In this paper, we introduce and discuss the framework of sampling using time-changed Markov processes to obtain more efficient continuous-time MCMC algorithms for challenging tasks, such as those described above. The fundamental idea is to apply a random time-change to a base Markov process in such a way that time accelerates in regions where we want the process to visit more frequently. At the same time, this ensures that the process does not spend too much time in these regions, which is crucial for the process to satisfy a law of large numbers. The interpretation of these regions depends on the specific task at hand. For instance, they may correspond to low-density areas that separate modes in a multimodal target setting or to the tails of the target distribution in the case of rare events estimation. In practice, the time-change is regulated by a speed function, $s$, which maps the state space to strictly positive real numbers and accelerates or decelerates time based on the current state of the process.
@@
Throughout the paper, we assume that the base process, denoted by $Y$, is a well-studied Markov process. We provide conditions on $Y$ and $s$ that ensure desirable properties of the time-changed process $X$, such as the law of large numbers (LLN), qualitative convergence rates to the stationary distribution, and functional central limit theorems. In particular, suitable choices of $s$ ensure that the convergence of $X$ is exponential in time, even if $Y$ converges only polynomially or more slowly. Moreover, we provide conditions for the uniform ergodicity of $X$, which refers to the exponential convergence of its law to the target, independently of the initial condition. Throughout the paper, we highlight the numerous connections to the existing literature and apply our framework to several Markov processes that are widely used in the context of MCMC algorithms.
Related work
The idea of sampling from a measure $\mu$ with a time-change of a Markov process was first discussed by Roberts and Stramer (2002a). The authors consider a time-changed overdamped Langevin diffusion corresponding to speed functions of the particular form $s(x)=\mu^{-\alpha}(x)$ for $\alpha\in(0,1)$, where $\mu$ is the target distribution. The focus of Roberts and Stramer (2002a) is then limited to a specific choice of process and speed function, whereas in this paper we define and study the framework in generality. The time-change of the Zig-Zag process (ZZP), a particular PDMP, was already considered by Vasdekis and Roberts (2023), but the process was introduced with ad-hoc arguments and without indicating this connection. The approach of time-change is also deeply connected to importance sampling (Robert and Casella, 2004; Owen, 2013). Indeed, computing ergodic averages with respect to a time-changed process is equivalent to obtaining re-weighted ergodic averages of the base process for a random time horizon (see Equation (8) below). In this direction, Chak et al. (2023) studies the optimal choice of “proposal distribution” $\tilde{\mu}$ to use as target for the overdamped Langevin diffusion, then adjusting the obtained sample by importance weights. We shall revisit this connection in Example 3.8. It is worth mentioning that the speed function $s(x)=\mu^{-\alpha}(x)$ with $\alpha=1$ was shown to be optimal across all diffusion coefficients for the overdamped Langevin diffusion in a specific (multimodal) setting when the state space is included in a compact set (Lelièvre et al., 2024). Time-changed Langevin diffusions were also considered by Leroy et al. (2024), who established some first theoretical properties of such processes and proposed suitable discretisation schemes for their practical simulation. In the discrete-time world, Andral et al. (2024) use as base process a Markov chain with stationary distribution different than the target and adjust it by letting it spend a random number of iterations at each state. This is closely connected to some of the ideas we discuss in Section 2.2. Similar ideas were introduced by Malefaki and Iliopoulos (2008) in a semi-Markov implementation. The frameworks of time changes and space transformations are similar when the main goal is to increase the frequency of visits to the tails, or equivalently to obtain a uniformly ergodic process. This aspect is discussed in Section 5, where we particularly focus on the connection to Johnson and Geyer (2012) and Yang et al. (2024). Finally, we note that a first version of this work appeared in the PhD thesis of the first author (Bertazzi, 2023).
@@
  • Section 5 discusses the relation between time changes and space transformations, illustrating the similarities and differences of the two methods. In particular, we shall observe that the speed function in our framework plays a similar role as the Jacobian determinant of the diffeomorphism that defines the space transformations. We analyse this connection considering a novel diffeomorphism, which maps $\mathbb{R}^{d}$ to points inside the $d$-dimensional unit sphere.
@@
  • LABEL:{sec:discussion} contains a brief discussion, summarizing the results of this paper and suggesting potential directions for future work.
  • The appendices contain the proofs of our theoretical results, along with additional numerical simulations.
@@ 2.2. Estimating expectations with time-changed Markov processes
    In cases where simulating the process $X$ is challenging, we can estimate $\mu(f)$ by re-weighting paths of the base process $Y$. Indeed, following the proof of Theorem 2.3 (see Appendix A), the change of variables $u=r(t)$, with Jacobian $s(Y_{u})$, gives
    $\frac{1}{T}\int_{0}^{T}f(X_{t})\textnormal{d}t\,=\,\frac{1}{T}\int_{0}^{r(T)}\frac{f(Y_{u})}{s(Y_{u})}\textnormal{d}u.$ (8)
    Theorem 2.3 guarantees that the right hand side of (8) is a consistent estimator of $\mu(f)$. The estimator (8) requires the simulation of $Y$ for a random time $r(T)$, which is not always streightforward to calculate. Alternatively, we may fix a deterministic time horizon $T$ and use the estimator
@@
    which is equivalent to simulating $X$ until time $r^{-1}(T)$. For the right hand side of (9) we used the expression given by (3)for the quantity $r^{-1}(T)$.
    The estimator (9) essentially coincides with doing self-normalised importance sampling for the base process $Y$, where the unnormalised weights are $(s(Y_{t}))^{-1}$. The estimator (9) is considered in Chak et al. (2023) for a particular choice of the base process, $Y$. Similarly to importance sampling, this approach allows the estimation of the normalising constant of $\mu$. Indeed, let $\mu(x)=Z^{-1}\exp(-\mathrm{U}(x))$, where $Z$ is unknown, but where we have access to the function $\mathrm{U}$. As will be seen in the proof of Theorem 2.3 (see Appendix A), we can estimate $Z$ using
    $\frac{r^{-1}(T)}{T\int_{\mathrm{E}}s(x)\exp(-\mathrm{U}(x))\textnormal{d}x}\,\,\xrightarrow[a.s.]{T\rightarrow\infty}\,\,Z.$ (10)
    This can be used when one has access to the quantity $\int_{\mathrm{E}}s(x)\exp(-\mathrm{U}(x))\textnormal{d}x$.
@@
    We now provide a third option in the case where directly simulating $X$ or the random time-change $r^{-1}(T)$ is not feasible. An alternative strategy is to use as base process $Y$ a jump process with jump rate $\tilde{\lambda}(y)=1$ for all $y\in\mathrm{E}$, and jump kernel $\widetilde{Q}$, that has $\tilde{\mu}$ as stationary distribution. The kernel $\widetilde{Q}$ can be any kernel from the literature on discrete time MCMC algorithms, or the transition kernel of a continuous time Markov process after some fixed time $\bar{t}$. In either case, the time changed process $X_{t}=Y_{r(t)}$ is a jump process with rate $\lambda(x)=s(x)$ and kernel $Q=\widetilde{Q}$ (see Theorem 3.1 below), with stationary distribution $\mu$. Fixing a time horizon $T$, and a starting position $x$, we can directly simulate the time-changed process $X$ using the procedure described in Algorithm 1.
    Input: Time horizon $T$, starting position $x$, speed function $s$, target $\mu$, a Markov transition kernel $\widetilde{Q}$ that leaves $\tilde{\mu}$ as in eq. 4 invariant
    Output: Jump process $X$ targeting $\mu$.
    Set $n\leftarrow 0$, $T_{0}=0$, $X_{0}=x$;
    while $T_{n}<T$ do
           Simulate $\tau_{n+1}\sim\mathrm{Exp}(s(X_{T_{n}}))$. Set the next jumping time to be $T_{n+1}=T_{n}+\tau_{n+1}$;
           For all $t\in\left[T_{n},\min\{T,T_{n+1}\}\right)$ set $X_{t}=X_{T_{n}}$;
           Simulate the next jump of the process $X_{T_{n+1}}\sim\widetilde{Q}(X_{T_{n}},\cdot)$;
    return $\left(X_{t}\right)_{t\in[0,T]}$.
    Algorithm 1 Time-changed Markov jump process implementation
    This process is straightforward to simulate and has stationary distribution $\mu$ under the conditions of Theorem 2.3. This procedure suggests the following estimator
    $\frac{1}{T}\int_{0}^{T}f(X_{t})\textnormal{d}t\,=\,\frac{1}{T}\sum_{n=0}^{N_{T}}f(X_{T_{n}})(T_{n+1}-T_{n})\,,$ (11)
    where $N_{T}$ is the time of the last jumping event before time $T$. One can “regularise” the estimator (11) by instead discretising the path $(X_{t})_{t\in[0,T]}$, that is using
@@
Remark 2.5.
Equation (8) reveals a connection between sampling using time-changed processes and importance sampling. In importance sampling, samples from a measure $\tilde{\mu}$ are re-weighted by $\mu/\tilde{\mu}$ to approximate expectations with respect to $\mu$. Similarly, the paths of the time-changed process $X$ are obtained by re-weighting paths from the $\tilde{\mu}$-stationary base process, where $\tilde{\mu}$ is defined in (4). Here, the time change assigns importance weights, where large weights are assigned to states where the speed function is small, i.e. where the time-changed process moves slowly. The speed function, defined as $\tilde{\mu}/\mu$ up to a constant factor, can then be interpreted as the reciprocal of the importance weight. Intuitively, a well-chosen speed function defines a distribution $\tilde{\mu}$ that makes the task of interest easier to solve, e.g. connecting the modes of $\mu$ when it is multimodal. As in importance sampling, it is essential to choose speed functions bounded below by a strictly positive constant, ensuring that $\tilde{\mu}$ has heavier tails than $\mu$.
@@ 3.1. Time-changes of piecewise deterministic Markov processes (PDMPs)
3.1.1. Running example: the Zig-Zag process
@@
One-dimensional bimodal target
@@
3.1.2. Other piecewise deterministic Markov processes
Taking advantage of Theorem 3.1, we now obtain several novel PDMPs with stationary distribution $\mu$. Using the same ideas as in Section 3.1.1, it is possible to define time-changes of PDMPs such as those designed for subsampling (see e.g. Bierkens et al. (2019a)), or of other PDMPs such as Wu and Robert (2020); Bierkens et al. (2020).
@@ 3.2. Diffusion processes
Assuming this SDE is well posed, we find that the solution has the measure with density $\mu(x,v)\propto\exp(-\mathrm{U}(x)-(1/2)\lvert v\rvert^{2})$ as stationary distribution. Leroy et al. (2024) consider the well-posedness of this time-changed SDE (as well as that of Example 3.8) and propose discretisation schemes based on splitting schemes.